OpenStudy (anonymous):

Line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to line AB ?

OpenStudy (anonymous):

5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0 Options

OpenStudy (anonymous):

your answer would be -5x-3y=0 I just got my math teacher to help me with this problem and with his knowledge and mine we both was getting -5x-3y=0

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

did i earn my metal

OpenStudy (anonymous):

lol